OpenStudy (anonymous):

Find an equation of the circle that has center (-3,2) and passes through (4,-6) .

OpenStudy (anonymous):

(x -h)^2 + (y - k)^2 =r^2 ^equation of a circle. plug things in.

OpenStudy (anonymous):

(h,k) being the center

OpenStudy (azureilai):

equation of a circle is: \[(x-h)^{2}+(y-k)^{2}=r^{2}\] (h,k) is the center substitute (x,y) to find r rewrite equation

OpenStudy (anonymous):

i do that and come to sort(113)

OpenStudy (anonymous):

sqrt*

OpenStudy (anonymous):

r=sqrt113?

OpenStudy (azureilai):

yep

OpenStudy (anonymous):

and my equation would be (x--3)+(y-2)=113

OpenStudy (azureilai):

Don't forget the squared part in the left side (Refer to the standard equation I wrote above) you can leave 113 as it is though since it is already squared.

OpenStudy (anonymous):

i see, thanks team

OpenStudy (anonymous):

(4+3)^(2)+(-6-2)^(2)=r^2 49+64=r^2 113=r^2 equation would be (x+3)^(2)+(y-2)^(2)=113 so yes you had it right:)

OpenStudy (anonymous):

woot thanks! <3